This is the word problem I am having problems with.
You would like to have $4000 in four years for a special vacation following graduation by making deposits at the end of every 6 months in an annuity that pays 7% compounded semiannually.
A. How much should you deposit at the end of every six months?
B. How much of the $4000 comes from deposits and how much comes from interest?
I have tried the formula P=A(r/n)/(1+r/n)^nt -1 but it is not coming out to the correct answers.

Answer by stanbon(75887)

You would like to have $4000 in four years for a special vacation following graduation by making deposits at the end of every 6 months in an annuity that pays 7% compounded semiannually.
A. How much should you deposit at the end of every six months?
B. How much of the $4000 comes from deposits and how much comes from interest?
I have tried the formula P=A(r/n)/(1+r/n)^nt -1 but it is not coming out to the correct answers.
-------------------------
P=A(r/n)/(1+r/n)^nt -1
---
P = 4000(0.07/2)(1 + (0.07/2)^(2*4) -1
---
P = 140(1.3168)-1
P = $183.35
=================
Cheers,
Stan H.